# LaCour!: Enabling Research on Argumentation in Hearings of the European Court of Human Rights

### v1

@@ Appendix B Details on preliminary experiments
    The following training setups were considered for our preliminary experiments. The full dataset of questions and opinions is also available on https://github.com/trusthlt/lacour-questions-opinions and contains all 2423 pairs of questions and opinions (or lack thereof) with a label for the opinion type and a boolean value for the presence of a question and an opinion.
    The first setup (Table 5) considers all questions and uses logistic regression to classify a) whether there was an opinion or not and b) the type of opinion. The questions are transformed into sentence embeddings with Sentence-BERT (all-MiniLM-L6v2).
    model LR MLP
    train size 304 304
    dev size 76 76
    test size 95 95
    n layers 1 5
    input size 384 384
    batch size {64, 32, 16} {64, 32, 16}
    optimizer Adam Adam
[…]
    random seed 5113 5113
    epochs 50 50
    random baseline (accuracy) [%] 43.86 43.86
    random baseline ($F_{1}$ score) [%] 35.15 35.15
    majority (all 0) baseline (accuracy) [%] 59.65 59.65
    majority (all 0) baseline ($F_{1}$ score) [%] 0.00 0.00
    model best (accuracy) [%] 59.65 59.65
    model best ($F_{1}$ score) [%] 0.00 0.00
    Table 5: Experimental setup and training parameters for logistic regression (LR) and multilayer perceptron (MLP)
    Due to the simple setup of these experiments, a negative result was expected. These experiments were also conducted to see if there is an artifact or leakage of the label in the question.
    The second setup (Table 6) performs further training on a pre-trained sentence transformer. In this case, the pre-trained model ‘all-MiniLM-L6-v2’ is used as a base model and trained in a contrastive manner. A sentence transformer ingests a pair of two sentences and outputs their cosine similarity. The idea is that similar sentences have a high cosine similarity. In our case, the training objective is that pairs of questions which are followed up by an opinion should be closer to each other, while question pairs with unequal labels should be further away from each other. We used a simple contrastive loss for this and constructed similar pairs based on whether the question appears in the same hearing and carries the same binary opinion label (whether a question was followed up with an opinion or not). Negative examples are question pairs which appear in the same hearing but do not have the same opinion label. Additional ‘soft’ samples could be generated by also considering questions from different hearings with equal or unequal labels. The trained model was evaluated on a test set with a binary classification evaluator which only considers unseen hard negative and positive samples.
    sample generation ‘hard’ samples ‘soft’ samples
    train size 942 84302
    evaluation size 182 182
    loss contrastive constrastive
    tokenizer roberta-base roberta-base
    batch size 64 64
    random seed 5113 5113
    epochs 50 10
    random baseline (accuracy) [%] 43.41 43.41
    random baseline ($F_{1}$ score) [%] 56.44 56.44
    majority baseline (accuracy) [%] 61.54 61.54
    majority baseline ($F_{1}$ score) [%] 76.19 76.19
    model before training (accuracy) [%] 62.64 62.64
    model before training ($F_{1}$ score) [%] 76.71 76.71
    model best (accuracy) [%] 61.54 60.99
    model best ($F_{1}$ score) [%] 75.86 75.77
    Table 6: Experimental setup for training a sentence transformer model initialized from all-MiniLM-L6-v2 with contrastive loss using question pairs as hard and soft samples
    html>

### v4

@@ Appendix B Experimental setup details
The following training setups were considered for our experiments. The full dataset of questions and opinions is also available on huggingface [^21] https://huggingface.co/datasets/TrustHLT/QandO and contains a partition with 475 pairs of questions and opinions (or lack thereof) with a label for the opinion type for classification. The second partition consists of 792 non-duplicate triplets of a question, a positive example and a negative example used for training a sentence-transformer model with MultipleNegativesRankingLoss. [^22] https://sbert.net/docs/package_reference/sentence_transformer/losses.html
For all classification models, except Llama-3, we train the base models on the training data for 20 epochs with a batch size of 8 and an adaptive learning rate of 2e-6. The classification model using Llama-3 8b was done with 4bit quantization and a LoRA configuration to reduce the computational requirements for training a Llama model. Training for 3 epochs already lead to overfitting, so we reduced the training to only run for 1 epoch. All configurations use class weights to account for the imbalanced training data. We run all models on 3 different random seeds (75, 953, 5113), average across them and monitor F1-score, balanced accuracy as well as the confusion matrix.
For the sentence-transformer model we use Legal-BERT as a base model and train for 150 epochs with a batch size of 32. The chosen seed is 5113 and the dataset is split to reflect the class distribution into 70% for training, 15% for validation and the remaining 15% for testing.
We make the code for all experiments available on github. [^23] https://github.com/trusthlt/lacour-questions-opinions